# Welfare and Cost Aggregation for Multi-Agent Control: When to Choose Which Social Cost Function, and Why?

### v2

@@ III-A Comparability Levels
Let $C:\mathbb{R}^{n}\to\mathbb{R}$ be a social cost function. A vector of transformations $\{\phi_{i}\in\Phi:i\in\mathcal{N}\}$ is called an invariance transformation under $C$ if, for any cost profile $\mathbf{J}=(J_{1},\dots,J_{n})$, the transformed profile
@@
is equivalent to $\mathbf{J}$ under $C$, meaning that for all $x,y\in\mathcal{X}$,
$C(\mathbf{J}(x))\leq C(\mathbf{J}(y))\quad\Longleftrightarrow\quad C(\mathbf{J}^{\prime}(x))\leq C(\mathbf{J}^{\prime}(y)).$
We present four levels of interpersonal comparability, defined via the corresponding invariance transformations. Notice that by imposing a specific invariance condition on the SCF $C$, we are equivalently defining an information filter that specifies what features of the individual cost profiles are relevant for the social ordering. For each comparability level, we comment on what kind of equivalence they induce on the costs $J_{i}$ in terms of measurability and comparability.
@@
The costs $J_{i}$ are determined up to any common strictly increasing transformation $\phi_{\text{\tiny OLC}}$:
@@
The costs $J_{i}$ are determined up to distinct positive affine transformations:
@@
The costs $J_{i}$ are determined up to any common scale factor $a$, but may have distinct offsets $b_{i}$:
@@
The costs $J_{i}$ are determined up to common positive affine transformations:
@@
Figure 2: Venn diagram of the different interpersonal comparability levels. Cardinal Full Comparability (CFC) is often unwittingly assumed, with unintended fairness consequences.
III-B Permissible Social Cost Functions
@@
Figure 3: Guide to the results of Sections II, III-A, and III-B.
Classical results in the literature — most notably by Sen [45, 37], d’Aspremont and Gevers [36], and Roberts [35] — demonstrate that if a Social Cost Functional (SCFL) satisfies (P) and (IIA), then its form is uniquely determined by the level of interpersonal comparability assumed. We first provide a technical result informally stated in [46], then formalize the SCF choice in Theorem 1, summarizing results from the literature above in the formalism of our paper (see Figure 3 for a visual guide).
@@
Let $\varepsilon=(\varepsilon_{1},\dots,\varepsilon_{n})\in\mathbb{R}_{++}^{n}$. Since, under (OLC), (CNC), (CUC), and (CFC), $\mathfrak{F}$ is invariant under constant shifts $J_{i}(x)\to J_{i}(x)+\alpha_{i}$ (with $\alpha_{i}$ independent of $x$ and $\varepsilon^{\prime}=\max_{i}\varepsilon_{i}$ for (CFC)), choose $\alpha_{i}=\varepsilon_{i}$ for all $i$ and define
$\tilde{J}_{i}(x)=J_{i}(x)+\varepsilon_{i}\quad\forall x\in\mathcal{X}.$
Then, for any $x,y\in\mathcal{X}$ with $x\succ_{\mathbf{J}}y$, invariance yields $x\succ_{\tilde{\mathbf{J}}}y$ and, since the shift is uniform in $x$, the (PC) condition holds with $\varepsilon^{\prime}=\varepsilon$. ∎
Theorem 1 (Choice of Social Cost Functions [35], [36]).
@@
The SCFL $\mathfrak{F}$ is represented [^1] An additional mild equity argument is needed, see [36]. by the SCF
@@
with $g:\mathbb{R}^{n}\to\mathbb{R}$ homogeneous of degree 1.333gg is homogeneous of degree 1 if $g(\lambda x)=\lambda\,g(x)$ for all $\lambda>0$. For example one could choose $\gamma\max(\cdot)$ with $0\leq\gamma\leq 1$, thus reflecting a balance between efficiency and equity.
Additionally, if one requires the axiom of Anonymity (A), then all the $c_{i}$ become equal, i.e. $c_{i}=c_{j},\forall i,j\in\mathcal{N}$, and the function $g$ must be invariant to permutations of the agents.
Comparability OLC CNC CUC CFC
SCF $C(x)$ $\max_{i}J_{i}(x)$
$\prod_{i}\left[J_{i}(x_{0})-J_{i}(x)\right]^{c_{i}}$
with benchmark $x_{0}$
$\prod_{i}\left[-J_{i}(x)\right]^{c_{i}}$
with $b_{i}=0$
and negative costs
$\sum_{i}c_{i}\,J_{i}(x)$
$\frac{1}{n}\sum_{i}J_{i}(x)\;+\;g\left(\left[\begin{smallmatrix}J_{1}(x)-\frac{1}{n}\sum_{i}{J}(x)\\ \vdots\\ J_{n}(x)-\frac{1}{n}\sum_{i}{J}(x)\end{smallmatrix}\right]\right)$
$g$ homogeneous of degree 1
Allowed operations on the SCF $C(x)$
TABLE I: Synoptic table representing the appropriate SCF and SCF operations for each comparability level.
III-C Admissible Operations on Social Cost Functions
@@
One needs to be particularly careful about using SCFs for these purposes because Lemma 1 only certifies their use for selecting the best social choice. It is allowed, however, to perform other operations on the values of SCF for different outcomes, as long as the result of such operations is invariant with respect to the transformations listed in Section III-A, which guarantees that the result of such operations is meaningful. This is demonstrated in Proposition 1.
@@
or, in case of (CNC), of the ratios $\dfrac{C(\mathbf{J}(x))}{C(\mathbf{J}(y))},\ x,y\in\mathcal{X}^{\prime}$.
@@
Can be found in [47]. ∎
In practical settings, one is concerned with the meaningful statements that can be made about the values of the SCF $C$ evaluated for different outcomes. For example, by selecting $z=w=v$ in (1), the ratios-of-differences takes the form
$t:=\frac{C(\mathbf{J}(x))-C(\mathbf{J}(v))}{C(\mathbf{J}(y))-C(\mathbf{J}(v))},$
@@
It should not be surprising that other operations on the SCF $C$ are not generally permissible: $C$ is a cardinal representation of the order imposed by the SCFL $\mathfrak{F}$. Therefore, additional measures, such as setting a fixed reference point or a fixed scale, are needed to make other operations (e.g., absolute differences or ratios) meaningful.
IV Examples
@@
IV-A Water Allocation
@@
Agricultural irrigation accounts for 70% of global freshwater use already, and projections are that irrigation-based food production will need to grow another 50% by 2050 due to climate change in combination with population growth [48]. As groundwater reservoirs deplete [49], water will need to be used more efficiently and prudently to avoid resource collapse [50], and allocation/priority rules are needed to avoid a “tragedy of the commons” [51]. Recent works have explored advanced control techniques for water allocation [52, 53, 54, 55].
@@
In irrigation systems, the value a farmer draws from a certain amount of allocated water is difficult to quantify, especially as agricultural farmland and resulting needs are heterogeneous due to crops requiring different amounts of water per hectare and across seasons, with all of it being precipitation dependent [48, 50]. Thus, one can only conclude that farmers generally experience increased benefits with higher water allocations, though comparing it across heterogeneous farmers is not appropriate [56] (Figure 4). As stated in [57], regulators are interested in designing schemes that robustly satisfy social welfare and justice criteria despite this unmeasurable heterogeneity. These considerations, in the context of comparability, correspond to Cardinal Non Comparability (CNC) of farmers’ costs.
@@
In such a case, the SCF becomes $\prod_{i}\left[\,q_{i}x^{(i)}\right]^{c_{i}}$. The solution of (2) then satisfies proportional fairness [58], which coincides with the proportional allocation
@@
(see derivation in the appendix of [47]).
Proportional allocation is a commonly used allocation procedure implemented in constituencies worldwide [59, 60, 56, 61]. Thus, using social choice theory arguments and analyzing the underlying comparability notion we were able to give a different perspective on the commonly used proportional allocation rule in water irrigation systems.
IV-B Traffic Control
@@
Figure 5: Traffic routing example illustrating different comparability assumptions and the associated optimal outcomes.
Fig. 5 shows an example of the commonly studied traffic routing problem [14, 15, 16, 17] with $40$ long-distance commuters travelling from $O_{1}$ through $O_{2}$ to $D$, and $40$ short-distance commuters travelling from $O_{2}$ to $D$ only. The link delays are as shown in Fig. 5(a). We next discuss whether travel delay costs of both commuter types can be compared and illustrate the consequences of different comparability assumptions.
@@
The prevalence of the sum of costs as SCF in the literature suggests that either (CUC) or (CFC) is implicitly assumed. (CUC) is often assumed in conjunction with tolling solutions [62], in which the unit of comparison is the monetary value of time [10], and the commuters’ dispensable incomes, i.e., the affine offsets $b_{i}$, are not considered in the cost functions. Works that instead explicitly account for income differences[24] can be classified under (CFC). (CFC) is also natural when considering the travel delays themselves as costs (and not how they convert to money), with the justification that everyone has 24 hours in a day.
@@
In addition to basic axioms (P), (IIA), Anonymity (A) is desired if commuters are to have equal access rights to roads. Therefore, under (CUC) the only appropriate SCF is $\sum_{i}J_{i}$, while a larger family of SCFs is permissible under (CFC) (see Table I). To balance a trade-off between total delays and delays of the worst off, one can choose SCF $\sum_{i}J_{i}+\gamma\max J_{i}$. In our example, $\gamma$ essentially dictates to what extent long-distance commuters should be compensated in $O_{2}\rightarrow D$ for their additional delay in $O_{1}\rightarrow O_{2}$. Under (CNC), since travel delay costs are nonnegative, one must resort to Partial Independence (PI), and choose a suitable reference outcome $x_{0}$. A natural choice of $x_{0}$ is the non-controlled traffic equilibrium, since it represents the status quo or disagreement point if negotiations to adopt new control policies fail, and is commonly adopted in transportation to certify that policies are Pareto improving [63, 27]. This leads to SCF $\prod_{i}\left(J_{i}^{\textup{eq}}-J_{i}\right)$, with $J_{i}^{\textup{eq}}$ denoting the equilibrium delay to commuter $i$.
@@
The SCF can be used in an optimization formulation to determine the socially optimal traffic outcome. Figs. 5(b)5(d) illustrate the results in our example network, indicating the optimal link delays and associated allocations of long-distance (blue) and short-distance (orange) commuters [^4] The allocations are to be interpreted in a frequentist sense, e.g., in Fig. 5(d) each commuter uses the fast/slow link half of the times.. Notice that under (CUC), it does not matter which commuter uses the fast/slow link (indicated black in Fig. 5(b)). While these different allocations may appear intuitively more or less fair or efficient, we emphasize that the appropriate notion of ‘fairness and efficiency’ is underpinned by the assumed and justifiable level of comparability.
IV-C Energy Curtailment
Electrical power distribution grids host an ever-increasing amount of renewable power generation (e.g., residential solar panels). These grids have finite power transfer capacity, dictated by the physical limits of the infrastructure, which is expensive and sometimes impossible to reinforce. Therefore, access needs to be regulated: when overproduction occurs and power needs to be exported from these generators to the rest of the grid, some of the generation needs to be curtailed (up to 10% of the power generated by new installations [64]).
In mathematical terms, for a pool of $N$ generators, the grid operators need to decide the power curtailments $\{x_{i}\}_{i=1}^{N}$ for each generator $i$. In deciding that, they need to ensure that the grid operational constraints (voltage and line current limits) are satisfied when generators produce $p_{i}-x_{i}$, where $p_{i}$ is their current potential production. Such a feasibility problem leaves the grid operators to decide what is the “best” curtailment: on the one hand, the transition towards a sustainable power system calls for maximizing the total generation from renewable sources (in some cases, by law [65]). On the other hand, the energy system is a shared infrastructure, and each generator expects fair access to it. See [66] for a review of fairness in energy systems and [67, 68, 69, 70, 71, 72] for concrete proposals for fair curtailment strategies. These works propose multiple solutions but provide little guidance about which one to select, except for an a posteriori quantitative assessment of various fairness indices, which are equally hard to choose and justify.

### v3

@@ 3.A Comparability Levels
Let $\mathfrak{F}:\mathcal{J}^{n}\to\mathcal{R}$ be a social cost functional. A vector of transformations $\{\phi_{i}\in\Phi:i\in\mathcal{N}\}$ is called an invariance transformation under $\mathfrak{F}$ if, for any cost profile $\mathbf{J}=(J_{1},\dots,J_{n})$, the transformed profile
@@
is equivalent to $\mathbf{J}$ under $\mathfrak{F}$, meaning that for all $x,y\in\mathcal{X}$,
$x\succsim_{\mathbf{J}}y\quad\Longleftrightarrow\quad x\succsim_{\mathbf{J}^{\prime}}y.$
We present four levels of interpersonal comparability, defined via the corresponding invariance transformations. Notice that by imposing a specific invariance condition, we are equivalently defining an information filter that specifies what features of the individual cost profiles are relevant for the social ordering. For each comparability level, we comment on what kind of equivalence they induce on the costs $J_{i}$ in terms of comparability.
@@
The costs are determined up to any common strictly increasing transformation $\phi_{\text{\tiny OLC}}$:
@@
The costs are determined up to distinct positive affine transformations:
@@
The costs are determined up to any common scale factor $a$, but may have distinct offsets $b_{i}$:
@@
The costs are determined up to common positive affine transformations:
@@
Figure 2: Venn diagram of the interpersonal comparability levels. Cardinal Full Comparability (CFC) is often unwittingly assumed, with unintended fairness consequences.
@@
3.B Permissible Social Cost Functions
Comparability OLC CNC CUC CFC
SCF $C(x)$ $\max\limits_{i}J_{i}(x)$
$\prod_{i}\left[J_{i}(x_{0})-J_{i}(x)\right]^{c_{i}}$
with benchmark $x_{0}$
$\prod_{i}\left[-J_{i}(x)\right]^{c_{i}}$
with $b_{i}=0$
and negative costs
$\sum\limits_{i}c_{i}\,J_{i}(x)$
$\frac{1}{n}\sum\limits_{i}J_{i}(x)+g\left(\left[\begin{smallmatrix}J_{1}(x)-\frac{1}{n}\sum\limits_{i}{J_{i}}(x)\\ \vdots\\ J_{n}(x)-\frac{1}{n}\sum\limits_{i}{J_{i}}(x)\end{smallmatrix}\right]\right)$
$g$ homogeneous of degree 1
Allowed operations on the SCF $C(x)$
Table 1: Synoptic table representing the appropriate SCF and SCF operations for each comparability level (Theorem 1).
Classical results in the literature — most notably by Sen [sen_interpersonal_1970, Sen1979_welfarism], d’Aspremont and Gevers [dAspremontGevers2002], and Roberts [roberts_interpersonal_1980] — demonstrate that if a Social Cost Functional (SCFL) satisfies (P) and (IIA), then its form is uniquely determined by the level of interpersonal comparability assumed. We first provide a technical result informally stated in [hammond2023roberts], then formalize the SCF choice in Theorem 1, summarizing results from the literature above in the formalism of our paper (see Figure 3 for a visual guide and Table 1 for a synoptic table).
@@ 3.B Permissible Social Cost Functions
Figure 3: Guide to the results of Sections 2 and 3.
@@
Let $\varepsilon=(\varepsilon_{1},\ldots,\varepsilon_{n})\in\mathbb{R}_{++}^{n}$ and set $\alpha=\min_{i}\varepsilon_{i}$. Define a common shift $\tilde{J}_{i}(x)=J_{i}(x)+\alpha$ for all $i$ and all $x\in\mathcal{X}$. Invariance under common additive shifts yields $x\succ_{\mathbf{J}}y\Rightarrow x\succ_{\tilde{\mathbf{J}}}y$ for every pair $x,y\in\mathcal{X}$. Let $\varepsilon^{\prime}=\alpha\mathbf{1}$. Then, component-wise, $\tilde{\mathbf{J}}(x)\geq\mathbf{J}(x)+\varepsilon^{\prime}$ and $\tilde{\mathbf{J}}(y)=\mathbf{J}(y)+\alpha\leq\mathbf{J}(y)+\varepsilon$. So the condition (PC) holds with $\varepsilon^{\prime}$, uniformly across profiles and pairs. ∎
Theorem 1 (Choice of Social Cost Functions [roberts_interpersonal_1980], [dAspremontGevers2002]).
@@
The SCFL $\mathfrak{F}$ is represented [^1] An additional mild equity argument is needed, see [dAspremontGevers2002]. by the SCF
@@
with $g:\mathbb{R}^{n}\to\mathbb{R}$ homogeneous of degree 1.333gg is homogeneous of degree 1 if $g(\lambda x)=\lambda\,g(x)$ for all $\lambda>0$. For example one could choose $\gamma\max(\cdot)$ with $0\leq\gamma\leq 1$, thus reflecting a balance between efficiency and equity, nicely axiomatized in [bossert].
Additionally, if one requires the axiom of Anonymity (A), then all the $c_{i}$ in (CNC) and (CUC) become equal, i.e. $c_{i}=c_{j},\forall i,j\in\mathcal{N}$, and the function $g$ in (CFC) must be invariant to permutations of the agents.
3.C Admissible Operations on Social Cost Functions
@@
One needs to be particularly careful about using SCFs for these purposes because Proposition 1 only certifies their use for selecting the best social choice. It is allowed, however, to perform other operations on the values of SCF for different outcomes, as long as the result of such operations is invariant with respect to the transformations listed in Section 3.A, which guarantees that the result of such operations is meaningful. This is stated in the following results.
Lemma 2.
Consider $n$ values $(w_{1},\dots,w_{n})\in\mathbb{R}^{n}$. Suppose $q(w_{1},\dots,w_{n})$ is a rational function satisfying, $\forall\,a\in\mathbb{R}_{++},\;\forall\,b\in\mathbb{R}$,
$q(w_{1},\dots,w_{n})\;=\;q(a\,w_{1}+b,\;\dots,\;a\,w_{n}+b)$
Then $q$ is a rational function of the ratios of differences:
$\frac{\,w_{i}-w_{j}\,}{\,w_{k}-w_{\ell}\,},\quad 1\leq i,j,k,\ell\leq n,\quad(w_{k}-w_{\ell}\neq 0).$
Conversely, any rational expression built from these ratios of differences is invariant under $t\mapsto a\,t+b$.
Proof.
As $q$ is a rational function, w.l.o.g we check the identity on the open region where $w_{1}>w_{n}$. In this region, choose the parameters $a=\frac{1}{w_{1}-w_{n}}>0,\quad b=-aw_{n}.$ Substituting these into the invariance condition yields
$q(w_{1},\dots,w_{n})\;=\;q\Bigl(1,\frac{w_{2}-w_{n}}{w_{1}-w_{n}},\dots,\frac{w_{n-1}-w_{n}}{w_{1}-w_{n}},0\Bigr).$
The right-hand side is a rational expression depending solely on the ratios $\frac{w_{k}-w_{n}}{w_{1}-w_{n}}$ (or, equivalently, any ratio of differences $\frac{w_{i}-w_{j}}{w_{k}-w_{\ell}}$). This holds on the entire domain, as the equality holds on an open set and $q$ is a rational function. The converse is immediate. ∎
@@
or, in case of (CNC), of the ratios $\dfrac{C(\mathbf{J}(x),\mathbf{J}(x_{0}))}{C(\mathbf{J}(y),\mathbf{J}(x_{0}))},\ x,y\in\mathcal{X}^{\prime}$ and $x_{0}$ being the reference point.
@@
We show that admissible transformations described in Section 3.A induce an affine group action on the SCF $C$. For (CUC) we obtain $C(\phi_{\text{CUC}}(\cdot))=aC(\cdot)+\sum_{i}c_{i}b_{i}$. For (CFC) we obtain $C(\phi_{\text{CFC}}(\cdot))=aC(\cdot)+nb$. For (CNC) we obtain $C(\phi_{\text{CNC}}(\cdot))=(\prod_{i}a_{i}^{c_{i}})C(\cdot).$ Thus, Lemma 2 can be applied to characterize the full set of rational functions that are invariant under such transformation, with the (CNC) reduced to ratios by picking $w_{j}=w_{k}$ and then dividing the ratio-of-differences by $w_{k}$. ∎
It should not be surprising that other operations on the SCF $C$ are not generally permissible: $C$ is a cardinal representation of the order imposed by the SCFL $\mathfrak{F}$. Therefore, additional measures, such as setting a fixed reference point or a fixed scale, are needed to make other operations (e.g., absolute differences or ratios) meaningful. For instance, by selecting $z=w=v$ in (1), the ratios-of-differences takes the form
$t:=\frac{C(\mathbf{J}(x))-C(\mathbf{J}(v))}{C(\mathbf{J}(y))-C(\mathbf{J}(v))},$ (2)
@@
As an example, consider the use of Price of Anarchy (PoA) to evaluate and optimize system performance [Chandan2024, Zhang2018, Piliouras2017, Wang2017, hill2023tradeoff]. PoA quantifies (in)efficiency of a mechanism via the ratio between the value that the SCF takes at the worst-case Nash equilibrium and at the social optimum. However, in its classical form, PoA is not invariant to the affine transformations and one needs to replace it with a metric that is based on the admissible operations defined in Proposition 2. One can adopt the ratio between the improvement of the Nash equilibrium and the improvement of the social optimum, both with respect to a benchmark solution (or, in the case of CNC, with respect to the reference point).
4 Examples
@@
4.A Water Allocation
Agricultural irrigation accounts for 70% of global freshwater use already, and projections are that irrigation-based food production will need to grow another 50% by 2050 due to climate change in combination with population growth [bwambale2022smart]. As groundwater reservoirs deplete [condon2021global], water will need to be used more efficiently and prudently to avoid resource collapse [li2017irrigation], and allocation/priority rules are needed to avoid a “tragedy of the commons” [olorunfemi2023how]. Recent works have explored advanced control techniques for water allocation [negenborn2009distributed, castelletti2023model, valledesma2024water, wang2025stochastic].
@@ 4.A Water Allocation
In irrigation systems, the value a farmer draws from a certain amount of allocated water is difficult to quantify, especially as agricultural farmland and resulting needs are heterogeneous due to crops requiring different amounts of water per hectare and across seasons, with all of it being precipitation dependent [bwambale2022smart, li2017irrigation]. Thus, one can only conclude that farmers generally experience increased benefits with higher water allocations, though comparing it across heterogeneous farmers is not appropriate [gomezlimon2020agricultural] (Figure 4). As stated in [madani2013exogenous], regulators are interested in designing schemes that robustly satisfy social welfare and justice criteria despite this unmeasurable heterogeneity. These considerations, in the context of comparability, correspond to Cardinal Non Comparability (CNC) of farmers’ costs.
@@
In such a case, the SCF becomes $\prod_{i}\left[\,q_{i}x^{(i)}\right]^{c_{i}}$. The solution of (3) then satisfies proportional fairness [Kelly1998], which coincides with the proportional allocation
@@
(see derivation in the appendix of [welfarism-arxiv]).
Proportional allocation is a commonly used allocation procedure implemented in constituencies worldwide [ostrom1994rules, roagarcia2014equity, gomezlimon2020agricultural, rosegrant1994markets]. Thus, using social choice theory arguments and analyzing the underlying notion of comparability, we can offer a different perspective on the commonly used proportional allocation rule in water irrigation systems.
4.B Traffic Control
Fig. 5 shows an example of the commonly studied traffic routing problem [Chandan2024, Zhang2018, Piliouras2017, Wang2017] with $40$ long-distance commuters travelling from $O_{1}$ through $O_{2}$ to $D$, and $40$ short-distance commuters travelling from $O_{2}$ to $D$ only. The link delays are as shown in Fig. 5(a). We next discuss whether travel delay costs of both commuter types can be compared and illustrate the consequences of different comparability assumptions.
@@ 4.B Traffic Control
Figure 5: Traffic example illustrating different comparability assumptions and the associated optimal outcomes.
@@
The prevalence of the sum of costs as SCF in the literature suggests that either (CUC) or (CFC) is implicitly assumed. (CUC) is often assumed in conjunction with tolling solutions [yang1998principle], in which the unit of comparison is the monetary value of time [zamparini2007meta], and the commuters’ dispensable incomes, i.e., the affine offsets $b_{i}$, are not considered in the cost functions. Works that instead explicitly account for income differences[jalota2021efficiency] can be classified under (CFC). (CFC) is also natural when considering the travel delays themselves as costs (and not how they convert to money), with the justification that everyone has 24 hours in a day.
@@
In addition to basic axioms (P), (IIA), Anonymity (A) is desired if commuters are to have equal access rights to roads. Therefore, under (CUC) the only appropriate SCF is $\sum_{i}J_{i}$, while a larger family of SCFs is permissible under (CFC) (see Table 1). To balance a trade-off between total delays and delays of the worst off, one can choose SCF $\sum_{i}J_{i}+\gamma\max J_{i}$. In our example, $\gamma$ essentially dictates to what extent long-distance commuters should be compensated in $O_{2}\rightarrow D$ for their additional delay in $O_{1}\rightarrow O_{2}$. Under (CNC), since travel delay costs are nonnegative, one must resort to Partial Independence (PI), and choose a suitable reference outcome $x_{0}$. A natural choice of $x_{0}$ is the non-controlled traffic equilibrium, since it represents the status quo or disagreement point if negotiations to adopt new control policies fail, and is commonly adopted in transportation to certify that policies are Pareto improving [daganzo2000pareto, elokda2024carma]. This leads to SCF $\prod_{i}\left(J_{i}^{\textup{eq}}-J_{i}\right)$, with $J_{i}^{\textup{eq}}$ denoting the equilibrium delay to commuter $i$.
@@
The SCF can be used in an optimization formulation to determine the socially optimal traffic outcome. Figs. 5(b)5(d) illustrate the results in our example network, indicating the optimal link delays and associated allocations of long-distance (blue) and short-distance (orange) commuters [^4] The allocations are to be interpreted in a frequentist sense, e.g., in Fig. 5(d) each commuter uses the fast/slow link half of the times.. Notice that under (CUC), it does not matter which commuter uses the fast/slow link (indicated black in Fig. 5(b)). While these different allocations may appear intuitively more or less fair or efficient, we emphasize that the appropriate notion of “fairness and efficiency” is underpinned by the assumed and justifiable level of comparability.
4.C Energy Curtailment
Electrical power distribution grids host an ever-increasing amount of renewable power generation (e.g., residential solar panels). These grids have finite power transfer capacity, dictated by the physical limits of the infrastructure, which is expensive and sometimes impossible to reinforce. Therefore, access needs to be regulated: when overproduction occurs and power needs to be exported from these generators to the rest of the grid, some of the generation needs to be curtailed (up to 10% of the power generated by new installations [NOVAN2024102930]).
In mathematical terms, for a pool of $N$ generators, the grid operators need to decide the power curtailments $\{x_{i}\}_{i=1}^{N}$ for each generator $i$. In deciding that, they need to ensure that the grid operational constraints (voltage and line current limits) are satisfied when generators produce $p_{i}-x_{i}$, where $p_{i}$ is their current potential production. Such a feasibility problem leaves the grid operators to decide what is the “best” curtailment: on the one hand, the transition towards a sustainable power system calls for maximizing the total generation from renewable sources (in some cases, by law [EEG2017]). On the other hand, the energy system is a shared infrastructure, and each generator expects fair access to it. See [soares2024review] for a review of fairness in energy systems and [DallAnese2014, gebbran2021fair, lusis2019reducing, liu2020fairness, Borbath2024, Moring2024Fair-Over-TimeCoordination] for concrete proposals for fair curtailment strategies. These works propose multiple solutions but provide little guidance about which one to select, except for an a posteriori quantitative assessment of various fairness indices, which are equally hard to choose and justify.